On a hot day, the temperature of a 33,000 L swimming pool increases by 1.70°C. What is the net heat transfer (in J) during this heating? Ignore any complications, such as loss of water by evaporation. (Enter the magnitude.)
Answer #1

mass of water, m = volume * density

m = 33000* 10^-3* 1000

m = 33000 kg

heat need to be transfered

H = m C ( dT)

H= 33000* 4186* 1.7

H = 2.348*10^8 J

=========

Comment in case any doubt, will reply for sure. Goodluck